formula of beat frequency

Grade:11

10 Answers

PRAPULPODISHETTI
8 years ago
Beat frequency=(1/Ta – 1/Tb).
KUNCHAM SAMPATH
8 years ago
Beat frequency=(1/Ta – 1/Tb).
L GOUSE BASHA
8 years ago
Beat frequency=(1/Ta – 1/Tb)
RAKESH CHINDAM
8 years ago
Beat frequency=(1/Ta – 1/Tb).
raju p
8 years ago
no of beats( slightly difference between two frequencies) per second is called BEAT FREQUENCY.....
raju p
8 years ago
THE FORMULA OF BEAT FREQUENCY=(1/Ta – 1/Tb)
Gavvala Ganesh
8 years ago
Beat frequency=(1/Ta – 1/Tb).
HARINATH
8 years ago
Beat frequency=(1/Ta – 1/Tb).
SEELA YUVARAJU
8 years ago
beat frequency=(1/Ta-Tb)
SAI SARDAR
8 years ago
the formula of beat frequency is =(1/Ta – 1/Tb).
here,
Ta = time period of first frequency.
Tb = time period of second frequency.